Using the following chemical equation how many grams of calcium chloride can be produced by reacting 57.09 g of calcium hydroxide?

2HCl + Ca(OH)2 -> 2H2O + CaCl2

1
Expert's answer
2021-02-15T02:47:44-0500

2HCl + Ca(OH)2 -> 2H2O + CaCl2

According to the equation, the mass of calcium chloride equals:

m(CaCl2) = m(Ca(OH)2) × Mr(CaCl2) / Mr(Ca(OH)2) = 57.09 g × 110.98 g/mol / 74.09 g/mol = 85.52 g


Answer: 85.52 g